I'm having a problem with putting together this word problem and I am not sure how to solve this. I am really confused. Can someone help?
A nurse must be careful not to inject a drug too quickly into a patient’s IV. If the drug is injected at the maximum rate of 0.03 cc per second, and the total amount to be injected is 15 cc,
a. Create an equation that tells the amount injected as a function of time (in seconds)
b. Determine the time required to complete the injection.

Answer by rfer(16322)

0.03/1=15/x
0.03x=15
x=15/0.03
x=500 sec